OpenStudy (dr0zier99):

please help will medel -3x+2y=2 i know that the slope is 3/2 and that the y intercept is 1 but i dont know how o graph it can someone help me how

OpenStudy (dr0zier99):

can you please help me

OpenStudy (er.mohd.amir):

put x=0,1,-1 give u y values

OpenStudy (dr0zier99):

ok i did 0,1

OpenStudy (dr0zier99):

what do you mean -1 give u y value

OpenStudy (er.mohd.amir):

whhen x=0 y=1 plote this point and put x=1 give y=5/2 and so on

OpenStudy (er.mohd.amir):

when x=-1 y=-1/2

OpenStudy (er.mohd.amir):

make a line pass all these points

OpenStudy (dr0zier99):

im sorry but im confused

OpenStudy (er.mohd.amir):

on coordinate axes (x,y) are (0,1), (1,5/2), (-1,-1/2) cross all these points on paper